can you help me solve this word problem? the word problem is Mr. Enigma said, "Five less than one third of my
age is 14." How old is Mr. Enigma?

Answer by Cromlix(4381)

1/3x - 5 = 14
1/3x = 14 + 5
1/3x = 19
x = 57
Mr Enigma is 57 years old.
Hope this helps.
:-)